OpenStudy (anonymous):

How do I calculate with De Moivre: z^4+36=0 ? z = complex number Thanks.

OpenStudy (arbershabani97):

z^4=36 z= \[\sqrt[4]{36}\]= 2.449489743

OpenStudy (arbershabani97):

oops it's -36 so the result is 2.449489743i

OpenStudy (anonymous):

There are 4 complex roots.. I need the trigonometric form :/ As x=cos(a)+isin(a)

OpenStudy (arbershabani97):

sorry but don't know anything about trigonometry

OpenStudy (anonymous):

thanks anyway :o Anyone? :( Medals given.